OpenStudy (anonymous):

I need help with this question. Rewrite the expression as a trigonometric function of a single angle measure. cos5xcosx +sin5xsinx

OpenStudy (vishweshshrimali5):

Okay see I have a formula: \(\large{\sin{(A+B)} = \sin{A}\cos{B} + cos{A}\sin{B}}\) Can you use this formula here ?

OpenStudy (vishweshshrimali5):

Here, A = 5x and B = x

OpenStudy (vishweshshrimali5):

@jhhj are you there ?